OpenStudy (anonymous):

what is the volume of a sphere with a diameter of 6 cm ( round to the nearest hundredth of a cm ) 4/3*3.41*33 = 122.76 cubic cm is this right?????

OpenStudy (lgbasallote):

you do mean \[\frac 43 \times 3.14 \times 3^3\]right?

OpenStudy (anonymous):

yes.

OpenStudy (lgbasallote):

then how did you get that answer?

OpenStudy (lgbasallote):

you should recheck your solution...what you got isnt right

OpenStudy (lgbasallote):

yes...i think you used 3.41...should be 3.14

OpenStudy (anonymous):

ok i will try it again

OpenStudy (anonymous):

113.04 ?????

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

thank you

OpenStudy (lgbasallote):

welcome ^_^